# On quantum learning algorithms for noisy linear problems

### v1

@@ Abstract
Quantum algorithms have shown successful results in solving noisy linear problems with quantum samples in which cryptographic hard problems are relevant. In this paper the previous results are investigated in detail, leading to new quantum and classical algorithms under the same assumptions as in the earlier works. To be specific, we present a polynomial-time quantum algorithm for solving the ring learning with errors problem with quantum samples which was deemed to be infeasible in [12], as well as polynomial-time classical algorithms that are more efficient than the corresponding quantum algorithms in solving the short integer solution problem with quantum samples and the learning with errors problem with size-reduced quantum samples.
@@
Quantum algorithm Bernstein-Vazirani algorithm Learning with errors Machine learning
1 Introduction
@@
In designing algorithms, setting underlying assumptions on what one can and cannot do is a crucial starting point. Indeed in a study of quantum algorithms, seemingly innocent assumptions sometimes become the center of controversies, for example one in a blackbox query [7, 9], in a state preparation [1], or on quantum random access memory (QRAM) [13]. Considering that the field is not in mere proof-of-concept stage, it is a necessary task to examine such assumption and evaluate its plausibility.
@@
One of the rapidly growing fields in quantum technology is machine learning where researchers try to make the best use of physical phenomena in studying the patterns in data. In 2019, an important milestone was achieved by Grilo et al. [12] where the authors have proven that the learning with errors (LWE) problem is efficiently solvable by using Bernstein-Vazirani (BV) algorithm [5], if a certain form of a superposed data sample is available. The paper also investigated the applicability of the learning algorithm to the learning parity with noise (LPN), the ring learning with errors (RLWE), the learning with rounding, and the short integer solution (SIS) problems, although not all were successful. The algorithm was further revised in 2022 by utilizing the divide-and-conquer strategy, basically dealing with component-wise problems [19].
@@
As clearly noted by the authors of [12], it does not directly mean the end of some lattice based cryptography, as it is not clear how such sample can be prepared. There has been an attempt to build a quantum circuit for QRAM in order to estimate resources for the sample preparation [20], but the very design requires exponentially many qubits which likely hinder the realization of the attack. On the opposite hand, there has been no mathematical proof that an efficient way of obtaining quantum samples defined in [12] is impossible. To sum up, it is too early to be optimistic (or pessimistic) on the plausibility of the learning algorithms along the line.
@@
Nevertheless, we have noticed in some occasions specific to [12, 19], certain assumptions are highly unlikely, or the negative prospect given by the authors turns out to be feasible. This work is summarized as follows:
  • Grilo et al. have argued in their paper (Appendix B in [12]) that their LWE algorithm cannot be extended to solve RLWE problems under the error model by Bshouty and Jackson [8]. After examining the issues and problems raised by the authors, we have come up with a solution leading to a quantum learning algorithm for the RLWE problem.
  • The same authors have claimed that the SIS problem can be solved by the learning algorithm with quantum samples. Here we show our concern that a quantum sample assumption in SIS is less plausible than that in LWE. In particular we show that given the SIS sample introduced by the authors, there exists a polynomial-time classical algorithm for the SIS problem.
@@
  • Divide-and-conquer algorithm proposed by [19] takes advantage of component-wise approach to the secret vector, suggesting size-reduced quantum samples. We question the practicality of such samples by proposing classical algorithms for the LWE problem given the same samples.
The rest of the paper is organized as follows. In Section 2, we define some notations and introduce hard problems relevant to noisy linear problems. A brief review of the Bernstein-Vazirani algorithm and how it brings an algorithm for solving the LWE problem with quantum samples are also covered. Section 3 introduces a new algorithm for solving the RLWE problem with quantum samples. An algorithm for the SIS problem with quantum samples devised by Grilo et al. [12] is reexamined in Section 4. In Section 5, we give polynomial-time classical algorithms for the LWE problem with size-reduced samples.
2 Preliminaries
@@
This section first introduces cryptographic hard problems related with noisy linear learning, and reviews the (generalized) Bernstein-Vazirani (BV) algorithm [5] followed by the algorithm of [12] for solving the LWE problem with quantum samples. Readers are assumed to be familiar with Bra-ket notation and the elementary quantum computation.
@@
    ${\bm{a}},{\bm{b}}\right>\;$ $=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{n-1}b_{n-1}.$
  • $\bullet$
@@
  • $\bullet$
    Let us denote $\mathsf{poly}(x_{1},\ldots,x_{m})=O(x_{1}^{c_{1}}\cdots x_{m}^{c_{m}})$ for some constants $c_{1},\ldots,c_{m}$.
2.1 Hard problems
@@
2.1.1 Learning parity with noise
Let ${\bm{s}}\in\{0,1\}^{n}$ be a secret vector and let $\eta\in[0,1/2)$ be a noise parameter. An LPN sample is given by
$({\bm{x}},\left<{\bm{s}},{\bm{x}}\right>+e\bmod 2)\in\{0,1\}^{n}\times\{0,1\}$
@@ 2.1.1 Learning parity with noise
2.1.2 Learning with errors
The LWE problem is a generalization of the LPN problem. Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector where $q\in\mathbb{Z}^{+}$, and let $\chi_{\textup{lwe}}$ be a probability distribution on $\mathbb{Z}$, called error distribution. An LWE sample is given by
$({\bm{a}},\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}$
where ${\bm{a}}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and $e\leftarrow\chi_{\textup{lwe}}$. The (search) LWE problem asks to find ${\bm{s}}$, given such LWE samples.
2.1.3 Ring learning with errors
The RLWE problem is similar to the LWE problem, but it deals with elements in the ring $\mathcal{R}:=\mathbb{Z}[X]/\langle\phi(X)\rangle$ instead of a vector in $\mathbb{Z}_{q}^{n}$, [^1] The RLWE problem is in general defined on a ring of integers of a number field, but for simplicity only polynomial rings are considered in this work. It is well-known that these two definitions are equivalent [18]. where $\phi(X)\in\mathbb{Z}[X]$ is a monic irreducible polynomial of degree $n$. Let $s(X)\in\mathcal{R}_{q}:=\mathcal{R}/q\mathcal{R}$ be a secret polynomial where $q\in\mathbb{Z}^{+}$. Let $\chi_{\textup{rlwe}}$ be a probability distribution on $\mathcal{R}$ which samples $a(X)=\sum_{i=0}^{n-1}a_{i}X^{i}\in\mathcal{R}$ as $a_{0},\ldots,a_{n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$. An RLWE sample is given by
$\big{(}a(X),s(X)\!\cdot\!a(X)+e(X)\bmod q\big{)}\in\mathcal{R}_{q}\times\mathcal{R}_{q}$
where $a(X)\leftarrow\mathcal{U}(\mathcal{R}_{q})$, $e(X)\leftarrow\chi_{\textup{rlwe}}$, and $\cdot$, $+$ denote multiplication and addition in $\mathcal{R}$, respectively. The (search) RLWE problem asks to find $s(X)$, given such RLWE samples.
2.1.4 Short Integer Solution
The SIS problem is defined as follows: given $n$ uniformly selected random vectors ${\bm{a}}_{1},\ldots,{\bm{a}}_{n}\in\mathbb{Z}_{q}^{m}$ for $q\in\mathbb{Z}^{+}$, find a nonzero ${\bm{v}}=(v_{1},\ldots,v_{n})\in\mathbb{Z}^{n}$ such that
$\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{0}}\bmod q\quad\textup{and}\quad\|{\bm{v}}\|\leq\beta$
@@ 2.1.4 Short Integer Solution
2.2 Bernstein-Vazirani algorithm
In this subsection, first we review how the BV algorithm works on binary vectors, and then generalize it.
2.2.1 Binary case
@@
Consider a vector ${\bm{s}}\in\{0,1\}^{n}$ and a function $f_{\bm{s}}:\{0,1\}^{n}\rightarrow\{0,1\}$, $f_{\bm{s}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod 2$. The problem to be tackled is
Given quantum oracle access to the unitary $U_{f_{\bm{s}}}:\lvert{\bm{x}}\rangle\lvert 0\rangle\mapsto\lvert{\bm{x}}\rangle\lvert f_{\bm{s}}({\bm{x}})\rangle$, find ${\bm{s}}$.
@@
Figure 1 (a) describes a quantum circuit for the BV algorithm [10, 5]. The initial state is $\lvert{\bm{0}}\rangle\lvert 1\rangle$, and then applying the gates as shown in Fig. 1 (a) achieves the goal. Details are not covered here, but the quantum state just before the measurement in Fig. 1 (a) reads,
$\frac{1}{2^{n}}\sum_{{\bm{y}}\in\{0,1\}^{n}}\sum_{{\bm{x}}\in\{0,1\}^{n}}(-1)^{\left<{\bm{s}}+{\bm{y}},{\bm{x}}\right>}\lvert{\bm{y}}\rangle\lvert-\rangle\enspace,$ (1)
@@
where $\lvert-\rangle=\frac{1}{\sqrt{2}}(\lvert 0\rangle-\lvert 1\rangle)$. As can be seen from Eq. (1), when measuring $\lvert{\bm{y}}\rangle$ register, the only non-trivial amplitude occurs for ${\bm{y}}={\bm{s}}$, and there does not exist any chance of measuring a string other than ${\bm{s}}$. Assuming single qubit gates are simultaneously executable and the query is efficient, the time complexity of the algorithm is $O(1)$.
Figure 1: (a) A circuit description of BV algorithm and (b) its variant. In (b), the Hadamard gate on the bottom wire and the subsequent measurement are not necessary, but included for easier analysis.
@@
Slightly different circuit is introduced in Fig. 1 (b). Using this circuit, the state before the measurement reads,
$\frac{1}{\sqrt{2}}\left(\lvert{\bm{0}}\rangle\lvert 0\rangle+\lvert{\bm{s}}\rangle\lvert 1\rangle\right)\enspace.$ (2)
@@
Notice that the Hadamard gate acting on the bottom wire is redundant, but only there to make Eq. (2) simple. It can be seen from Eq. (2) that measuring the first register gives ${\bm{s}}$ with probability 1/2. The procedure can simply be repeated until we get ${\bm{s}}$.
2.2.2 Generalized BV algorithm
@@
Instead of qubits, consider $q$-dimensional qudits. Each wire in Fig. 1 represents $q$-dimensional complex vectors and Hadamard gate is accordingly generalized [21]. Recall that generalized Hadamard operation can be understood as quantum Fourier transform on $\mathbb{Z}_{q}$[14] such as
@@
$\mathsf{QFT}_{\mathbb{Z}_{q}}:\lvert x\rangle$ $\mapsto$ $\sum_{y\in\mathbb{Z}_{q}}e^{2\pi\imath\frac{xy}{q}}\lvert y\rangle\enspace.$
@@
We are ready to describe the generalized BV algorithm. Notice that after the query in Fig. 1 (b), we would have $\frac{1}{\sqrt{q^{n}}}\sum_{{\bm{x}}\in\mathbb{Z}_{q}^{n}}\lvert{\bm{x}}\rangle\lvert f_{{\bm{s}}}({\bm{x}})\rangle$ where $f_{{\bm{s}}}:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{Z}_{q}$, $f_{{\bm{s}}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod q$. Instead of specifying a circuit for preparing such states, we simply define a quantum sample $\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$, and take it as an input to the BV algorithm as follows:
@@
$\lvert\psi\rangle\in(\mathbb{C}^{q})^{n+1}$
@@
3:
4:Apply $\mathsf{QFT}_{\mathbb{Z}_{q}}^{n+1}$ on $\lvert\psi\rangle$
5:Measure all qudits to get $({\bm{x}},y)\in\mathbb{Z}_{q}^{n}\!\times\!\mathbb{Z}_{q}$
6:if $\gcd(y,q)=1$ then
7:     return $\tilde{{\bm{s}}}:=-y^{-1}{\bm{x}}\bmod q$triangleright component-wise modular division
8:end if
9:return $\bot$
@@
One may think of the input $\lvert\psi\rangle$ as what we get by querying an oracle $\mathcal{O}_{\bm{s}}$. This algorithm will be the core subroutine in the learning algorithms.
@@
At this point, it will be a natural question to ask if the algorithm still works for noisy input samples. The main contribution of [12] is to answer the question as briefly reviewed below.
2.3 LWE algorithm with quantum samples
2.3.1 Quantum sample
@@
Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector and let $\mathcal{V}\subseteq\mathbb{Z}_{q}^{n}$. A quantum LWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$, is defined as follows:
@@
$\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ $=$ $\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{{\bm{a}}\in\mathcal{V}}\lvert{\bm{a}}\rangle\lvert\left<{\bm{s}},{\bm{a}}\right>+e_{\bm{a}}\bmod q\rangle,\quad e_{\bm{a}}\leftarrow\chi_{\textup{lwe}}$ (3)
@@
where $\chi_{\textup{lwe}}$ is a probability distribution on $\mathbb{Z}$. We assume that the support of $\chi_{\textup{lwe}}$ is $[-k,k]\cap\mathbb{Z}$ for some small positive integer $k$, i.e. $|e_{{\bm{a}}}|\leq k$. Define a quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ which outputs $\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle$ when queried.
2.3.2 Algorithm
Given access to the quantum oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$, one can find the secret vector ${\bm{s}}$ by using the $\mathsf{qLWE}\textup{-}\mathsf{Solver}$ described in Algorithm 2.
Algorithm 2 $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
1:oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
2:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$ or $\bot$
3:
4:$i\leftarrow 0$
5:repeat
6:     $\lvert\psi^{\textup{lwe}}_{{\bm{s}},V}\rangle\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright oracle query
7:     $\tilde{{\bm{s}}}\leftarrow\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$
8:     if  $\tilde{{\bm{s}}}=\bot$  then
9:         return $\bot$
10:     end if
11:     if  $\mathsf{Test}(\tilde{{\bm{s}}},\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}})=1$ then
12:         return $\tilde{{\bm{s}}}$
13:     end if
14:     $i\leftarrow i+1$
15:until $i\leq\ell_{1}$
16:return $\bot$
The subroutine $\mathsf{Test}$ in Algorithm 2 is designed to check if the candidate $\tilde{{\bm{s}}}$ is indeed the secret vector ${\bm{s}}$ by using additional LWE samples. Note that in Algorithm 3 the oracle $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$ can be quantum or classical.
Algorithm 3 $\mathsf{Test}$
1:$\tilde{{\bm{s}}}\in\mathbb{Z}_{q}^{n}$, oracle access to $\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$
2:$1$ (pass) or 00 (fail)
3:
4:$i\leftarrow 0$
5:repeat
6:     $({\bm{a}},b)\leftarrow\mathcal{O}^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}$triangleright $b=\left<{\bm{a}},{\bm{s}}\right>+e_{\bm{a}}\bmod q$
7:     if $|b-\left<{\bm{a}},\tilde{{\bm{s}}}\right>\bmod q|>k$ then
8:         return 00
9:     end if
10:     $i\leftarrow i+1$
11:until $i\leq\ell_{2}$
12:return $1$
2.3.3 Performance
Detailed analysis on Algorithm 2 is not covered here. Interested readers are encouraged to refer to Section IV B in [12]. Results and parameter choices are summarized below.
Success probability of $\mathsf{BV}_{q}$
Probability that $\mathsf{BV}_{q,n}(\lvert\psi^{\textup{lwe}}_{{\bm{s}},\mathcal{V}}\rangle)$ returns the correct secret ${\bm{s}}$ is lower bounded by
$\mathrm{\Omega}\!\left(\frac{|\mathcal{V}|}{kq^{n}}\right)\enspace.$
Pass/fail probability of $\mathsf{Test}$
If $\tilde{{\bm{s}}}={\bm{s}}$, $\mathsf{Test}$ returns $1$ (pass) with certainty. If $\tilde{{\bm{s}}}\neq{\bm{s}}$, the probability that $1$ is returned is at most
$\left(\frac{2k\!+\!1}{q}\right)^{\!\ell_{2}}\enspace.$
Parameter choices for $\mathsf{qLWE}\textup{-}\mathsf{Solver}$
By setting $\mathcal{V}=\mathbb{Z}_{q}^{n}$, $\ell_{1}=O(k\log(1/\eta))$, and $\ell_{2}=1$, we have

### v2

# Assessing the feasibility of quantum learning algorithms for noisy linear problems

@@ Abstract
Quantum algorithms for solving noisy linear problems are reexamined, under the same assumptions taken from the existing literature. The findings of this work include on the one hand extended applicability of the quantum Fourier transform to the ring learning with errors problem which has been left open by Grilo et al., who first devised a polynomial-time quantum algorithm for solving noisy linear problems with quantum samples. On the other hand, this paper also shows there exist efficient classical algorithms for short integer solution and size-reduced learning with errors problems if the quantum samples used by the previous studies are given.
@@
Quantum algorithm, Quantum Fourier Transform, Learning with errors, Machine learning, Bernstein-Vazirani algorithm
Introduction
@@
In designing algorithms, setting underlying assumptions on what one can and cannot do is a crucial starting point. Indeed in a study of quantum algorithms, seemingly innocent assumptions sometimes become the center of controversies, for example one in a blackbox query [1, 2], in a state preparation [3], or on quantum random access memory [4]. Considering that the field is not in the mere proof-of-concept stage, it is a necessary task to examine such an assumption and evaluate its plausibility.
@@
One of the rapidly growing fields in quantum technology is machine learning where researchers try to make the best use of physical phenomena in studying the patterns in data. Following the pioneering work on the quantum approach to learning parity with noise (LPN) problem by Cross et al. [5], an important milestone was achieved in 2019 by Grilo et al. [6] where the authors have proven that learning with errors (LWE) problem is efficiently solvable by using Bernstein-Vazirani (BV) algorithm [7], if a certain form of a superposed data sample is available. The paper also investigated the applicability of the learning algorithm to LPN, ring learning with errors (RLWE), learning with rounding, and short integer solution (SIS) problems, although not all were successful. The algorithm was further revised in 2022 by utilizing the divide-and-conquer strategy, basically dealing with component-wise problems [8].
@@
As clearly noted by the authors of the paper [6], it does not directly mean the end of some lattice based cryptography, as it is not clear how such a sample can be prepared. At the time of writing, there has been no known way to efficiently prepare such samples, but on the other hand, there has been no mathematical proof that an efficient way is nonexistent. To sum up, it is too early to be optimistic (or pessimistic) on the plausibility of the learning algorithms along the line.
@@
Nevertheless, we have noticed in some occasions specific to two papers [6, 8], certain assumptions allow polynomial-time classical algorithms questioning their soundness, or the negative prospect given by the authors turns out to be feasible. This work is summarized as follows:
  • The LWE algorithm with quantum samples is extended to solve RLWE problem, which has been deemed to be infeasible under the error model by Bshouty and Jackson [6, 9].
  • It is shown that the quantum sample assumption in the SIS algorithm is stronger than that in LWE in the sense that a polynomial-time classical algorithm can be devised given the same sample introduced by the previous work.
@@
  • Divide-and-conquer algorithm takes advantage of the component-wise approach to the secret vector, assuming size-reduced quantum samples [8]. We develop classical algorithms for the LWE problem given the same samples.
In the following, background information is covered prioritizing introduction to the LWE algorithm with quantum samples. New findings are then followed.
Preliminaries
@@
This section first introduces cryptographic hard problems related with noisy linear learning, and reviews the (generalized) Bernstein-Vazirani (BV) algorithm [7] for solving the LWE problem with quantum samples. Readers are assumed to be familiar with Bra-ket notation and the elementary quantum computation.
@@
    ${\bm{a}},{\bm{b}}\right>=a_{0}b_{0}+a_{1}b_{1}+\cdots+a_{n-1}b_{n-1}\enspace.$
@@
  • Let $\mathsf{poly}(x_{1},\ldots,x_{m})=O(x_{1}^{c_{1}}\cdots x_{m}^{c_{m}})$ for some constants $c_{1},\ldots,c_{m}$.
Hard problems
@@
Learning parity with noise. Let ${\bm{s}}\in\{0,1\}^{n}$ be a secret vector and let $\eta\in[0,1/2)$ be a noise parameter. An LPN sample is given by
$({\bm{x}},\left<{\bm{s}},{\bm{x}}\right>+e\bmod 2)\in\{0,1\}^{n}\times\{0,1\}\enspace,$
@@
Learning with errors. The LWE problem is a generalization of the LPN problem. Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector where $q\in\mathbb{Z}^{+}$, and let $\chi_{\textup{\tiny lwe}}$ be a probability distribution on $\mathbb{Z}$, called error distribution. An LWE sample is given by
$({\bm{a}},\left<{\bm{s}},{\bm{a}}\right>+e\bmod q)\in\mathbb{Z}_{q}^{n}\times\mathbb{Z}_{q}\enspace,$
where ${\bm{a}}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{n})$ and $e\leftarrow\chi_{\textup{\tiny lwe}}$. The (search) LWE problem asks to find ${\bm{s}}$, given such LWE samples.
Ring learning with errors. The RLWE problem is in general defined on a ring of integers of a number field, but for simplicity only polynomial rings are considered in this work. It is well-known that these two definitions are equivalent [16].
The RLWE problem is similar to the LWE problem, but it deals with elements in the ring $\mathcal{R}:=\mathbb{Z}[X]/\langle\phi(X)\rangle$ instead of a vector in $\mathbb{Z}_{q}^{n}$, where $\phi(X)\in\mathbb{Z}[X]$ is a monic irreducible polynomial of degree $n$. Let $s(X)\in\mathcal{R}_{q}:=\mathcal{R}/q\mathcal{R}$ be a secret polynomial where $q\in\mathbb{Z}^{+}$. Let $\chi_{\textup{\tiny rlwe}}$ be a probability distribution on $\mathcal{R}$ which samples $e(X)=\sum_{i=0}^{n-1}e_{i}X^{i}\in\mathcal{R}$ as $e_{0},\ldots,e_{n-1}\stackrel{{\scriptstyle i.i.d.}}{{\longleftarrow}}\chi$ using a probability distribution $\chi$ on $\mathbb{Z}$. An RLWE sample is given by
$\big{(}a(X),s(X)\!\cdot\!a(X)+e(X)\bmod q\big{)}\in\mathcal{R}_{q}\times\mathcal{R}_{q}\enspace,$
where $a(X)\leftarrow\mathcal{U}(\mathcal{R}_{q})$, $e(X)\leftarrow\chi_{\textup{\tiny rlwe}}$, and $\cdot$, $+$ denote multiplication and addition in $\mathcal{R}$, respectively. The (search) RLWE problem asks to find $s(X)$, given such RLWE samples.
Short integer solution. The SIS problem is defined as follows: given $n$ uniformly selected random vectors ${\bm{a}}_{1},\ldots,{\bm{a}}_{n}\in\mathbb{Z}_{q}^{m}$ for $q\in\mathbb{Z}^{+}$, find a nonzero ${\bm{v}}=(v_{1},\ldots,v_{n})\in\mathbb{Z}^{n}$ such that
$\sum_{i=1}^{n}v_{i}{\bm{a}}_{i}\equiv{\bm{0}}\bmod q\quad\textup{and}\quad\|{\bm{v}}\|\leq\beta\enspace,$
@@
Bernstein-Vazirani algorithm
Figure 1: (a) A circuit description of BV algorithm and (b) its variant
@@
Binary case. Consider a vector ${\bm{s}}\in\{0,1\}^{n}$ and a function $f_{\bm{s}}:\{0,1\}^{n}\rightarrow\{0,1\}$, $f_{\bm{s}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod 2$. The problem is, for given quantum oracle access to the unitary $U_{f_{\bm{s}}}:\left|{\bm{x}}\,\right\rangle\left|0\,\right\rangle\mapsto\left|{\bm{x}}\,\right\rangle\left|f_{\bm{s}}({\bm{x}})\,\right\rangle$, to find ${\bm{s}}$.
@@
Figure 1 (a) describes a quantum circuit for the BV algorithm [17, 7]. The initial state is $\left|{\bm{0}}\,\right\rangle\left|1\,\right\rangle$, and then applying the gates as shown in Figure 1 (a) achieves the goal. Details are not covered here, but the quantum state just before the measurement in Figure 1 (a) reads,
$\frac{1}{2^{n}}\sum_{{\bm{y}}\in\{0,1\}^{n}}\sum_{{\bm{x}}\in\{0,1\}^{n}}(-1)^{\left<{\bm{s}}+{\bm{y}},{\bm{x}}\right>}\left|{\bm{y}}\,\right\rangle\left|-\,\right\rangle\enspace,$ (1)
@@
where $\left|-\,\right\rangle=\frac{1}{\sqrt{2}}(\left|0\,\right\rangle-\left|1\,\right\rangle)$. As can be seen from Equation (1), when measuring $\left|{\bm{y}}\,\right\rangle$ register, the only non-trivial amplitude occurs for ${\bm{y}}={\bm{s}}$, and there does not exist any chance of measuring a string other than ${\bm{s}}$. Assuming single qubit gates are simultaneously executable and the query is efficient, the time complexity of the algorithm is $O(1)$.
@@
Slightly different circuit is introduced in Figure 1 (b). Using this circuit, the state before the measurement reads,
$\frac{1}{\sqrt{2}}\left(\left|{\bm{0}}\,\right\rangle\left|0\,\right\rangle+\left|{\bm{s}}\,\right\rangle\left|1\,\right\rangle\right)\enspace.$ (2)
@@
Notice that the Hadamard gate acting on the bottom wire is redundant, but only there to make Equation (2) simple. It can be seen from Equation (2) that measuring the first register gives ${\bm{s}}$ with probability 1/2. The procedure can simply be repeated until we get ${\bm{s}}$.
@@
Generalized BV algorithm. Instead of qubits, consider $q$-dimensional qudits. Each wire in Figure 1 represents $q$-dimensional complex vectors and Hadamard gate is accordingly generalized [18]. Recall that generalized Hadamard operation can be understood as quantum Fourier transform on $\mathbb{Z}_{q}$ [19] such as
@@
$\mathsf{QFT}_{\mathbb{Z}_{q}}:\left|x\,\right\rangle\mapsto\sum_{y\in\mathbb{Z}_{q}}e^{2\pi\imath\frac{xy}{q}}\left|y\,\right\rangle\enspace.$
@@
We are ready to describe the generalized BV algorithm. Notice that after the query in Figure 1 (b), we would have $\frac{1}{\sqrt{q^{n}}}\sum_{{\bm{x}}\in\mathbb{Z}_{q}^{n}}\left|{\bm{x}}\,\right\rangle\left|f_{{\bm{s}}}({\bm{x}})\,\right\rangle$ where $f_{{\bm{s}}}:\mathbb{Z}_{q}^{n}\rightarrow\mathbb{Z}_{q}$, $f_{{\bm{s}}}({\bm{x}})=\left<{\bm{s}},{\bm{x}}\right>\bmod q$. Instead of specifying a circuit for preparing such states, we simply define a quantum sample $\left|\psi\,\right\rangle\in(\mathbb{C}^{q})^{n+1}$, and take it as an input to the BV algorithm as follows:
@@
$\left|\psi\,\right\rangle\in(\mathbb{C}^{q})^{n+1}$
@@
3:Apply $\mathsf{QFT}_{\mathbb{Z}_{q}}^{n+1}$ on $\left|\psi\,\right\rangle$
4:Measure all qudits to get $({\bm{x}},y)\in\mathbb{Z}_{q}^{n}\!\times\!\mathbb{Z}_{q}$
5:if $\gcd(y,q)=1$ then
6:     return $\tilde{{\bm{s}}}:=-y^{-1}{\bm{x}}\bmod q$
7:end if
8:return $\bot$
@@
One may think of the input $\left|\psi\,\right\rangle$ as what we get by querying an oracle $\mathcal{O}_{\bm{s}}$. This algorithm will be the core subroutine in the learning algorithms.
@@
At this point, it will be a natural question to ask if the algorithm still works for noisy input samples. The main contribution of the previous work [6] is to answer the question as briefly reviewed below.
LWE algorithm with quantum samples
@@
Let ${\bm{s}}\in\mathbb{Z}_{q}^{n}$ be a secret vector and let $\mathcal{V}\subseteq\mathbb{Z}_{q}^{n}$. A quantum LWE sample for ${\bm{s}}$ and $\mathcal{V}$, denoted by $\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$, is defined as follows:
@@
$\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle=\frac{1}{\sqrt{|\mathcal{V}|}}\sum_{{\bm{a}}\in\mathcal{V}}\left|{\bm{a}}\,\right\rangle\left|\left<{\bm{s}},{\bm{a}}\right>+e_{\bm{a}}\bmod q\,\right\rangle,\leavevmode\nobreak\ e_{\bm{a}}\leftarrow\chi_{\textup{\tiny lwe}}\enspace,$ (3)
@@
where $\chi_{\textup{\tiny lwe}}$ is a probability distribution on $\mathbb{Z}$. We assume that the support of $\chi_{\textup{\tiny lwe}}$ is $[-k,k]\cap\mathbb{Z}$ for some small positive integer $k$, i.e. $|e_{{\bm{a}}}|\leq k$. Define a quantum oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$ which outputs $\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle$ when queried.
Given access to the quantum oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$, one can find the secret vector ${\bm{s}}$ by using the $\mathsf{qLWE}\textup{-}\mathsf{Solver}$ described in Algorithm 2 (see below). The subroutine $\mathsf{Test}$ in Algorithm 2 is designed to check if the candidate $\tilde{{\bm{s}}}$ is indeed the secret vector ${\bm{s}}$ by using additional LWE samples. Note that in Algorithm 3 the oracle $\mathcal{O}^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}$ can be quantum or classical. Detailed analysis on Algorithm 2 is not covered here. Interested readers are encouraged to refer to Section IV B in the previous paper [6]. Results and parameter choices are summarized as follows. Probability that $\mathsf{BV}_{q,n}(\displaystyle\left|\psi^{\textup{\tiny lwe}}_{{\bm{s}},\mathcal{V}}\,\right\rangle)$ returns the correct secret ${\bm{s}}$ is lower bounded by $\mathrm{\Omega}\!\left({|\mathcal{V}|}/({kq^{n}})\right)$. If $\tilde{{\bm{s}}}={\bm{s}}$, $\mathsf{Test}$ returns $1$ (pass) with certainty. If $\tilde{{\bm{s}}}\neq{\bm{s}}$, the probability that $1$ is returned is at most $\left(({2k\!+\!1})/{q}\right)^{\!\ell_{2}}$. By setting $\mathcal{V}=\mathbb{Z}_{q}^{n}$, $\ell_{1}=O(k\log(1/\eta))$, and $\ell_{2}=1$, we have